# Long-VITA: Scaling Large Multi-modal Models to 1 Million Tokens with Leading Short-Context Accuracy

### v2

@@ Abstract
We introduce Long-VITA, a simple yet effective large multi-modal model for long-context visual-language understanding tasks. It is adept at concurrently processing and analyzing modalities of image, video, and text over 4K frames or 1M tokens while delivering advanced performances on short-context multi-modal tasks. We propose an effective multi-modal training schema that starts with large language models and proceeds through vision-language alignment, general knowledge learning, and two sequential stages of long-sequence fine-tuning. We further implement context-parallelism distributed inference and logits-masked language modeling head to scale Long-VITA to infinitely long inputs of images and texts during model inference. Regarding training data, Long-VITA is built on a mix of $17$M samples from public datasets only and demonstrates the state-of-the-art performance on various multi-modal benchmarks, compared against recent cutting-edge models with internal data. Long-VITA is fully reproducible and supports both NPU and GPU platforms for training and testing. By leveraging our inference designs, Long-VITA models achieve a remarkable 2$\times$ prefill speedup and 4$\times$ context length extension in single node with $8$ GPUs. We hope Long-VITA can serve as a competitive baseline and offer valuable insights for the open-source community in advancing long-context multi-modal understanding.
@@ 1 Introduction
To further push the limits of open-source model capabilities, we extend the context length to $1$ million tokens and introduce Long-VITA, a strong open-source long-context visual language model. To this end, we employ a phased training approach that positions language as the pivot. Specifically, Long-VITA ’s ability to handle extended contexts is systematically augmented through a four-stage process. Beyond the conventional stages of vision-language alignment and supervised fine-tuning, our approach incorporates specialized stages for long-context supervised fine-tuning of $128$K and $1$M. To achieve a good performance trade-off between the long and short sequences, Long-VITA takes advantage of the existing abundance of open-source image-text and video-text data. We also introduce a multi-image summarization dataset, Comic-9K, comprising $9$k comic books and the corresponding detailed synopsis. This dataset has a total of $200$k images with an average of $20$ high-resolution photos for each sample, and the synopsis is all manually written collecting from the web.
@@
  • We evaluate Long-VITA on a wide range of benchmarks. Although Long-VITA is trained on open-source data only, comprehensive evaluations reveal that Long-VITA has emerged as the strong vision-language model among previous models of similar scale, especially in evaluating hallucinations and video understanding.
@@ 2.2 Long-Context Multi-Modal Model
Many methods have also been proposed in large multi-modal models to handle long-context visual inputs. LongVILA [11] execute a continuation of pre-training on the LLM to enhance its context length to $256$K, followed by long video training. LongLLaVA [12] integrates a hybrid of Mamba and Transformer blocks for long-context multi-modal understanding. LongVU [13] proposes a spatio-temporal adaptive compression scheme to reduce long video tokens by leveraging cross-modal query and inter-frame similarities. LongVA [14] extends the language model on text data and then aligns the extended model with visual inputs, which perceives more than $200$K visual tokens. Kangaroo [27] develops a curriculum training strategy that progressively equips the LLM basement with the capacity to comprehend long videos.
@@ 2.3 Long-Context Visual Instruction Data
LLaVA-Video [28] synthesizes long video-language instruction data, covering various tasks such as captioning, open-ended, and multi-choice QA. LongVILA [11] constructs instruction-following datasets from long videos, which encompasses summarization and other queries relevant to a comprehensive understanding of the content of long videos. Video-MME [29] incorporates a diverse range of video types, varying temporal durations, ranging from $11$ seconds to $1$ hour. LVBench [30] evaluates long-context LMMs that feature video-language interleaved inputs up to an hour long. LongVideoBench [31] introduces referring reasoning questions and presents significant challenges for both proprietary and open-source LMMs in their long-context multi-modal capabilities.
However, the above works only focus on long-context understanding in videos, which usually contain redundant visual frames and other modal information, such as subtitles and audio. In this paper, we explore the comic-based long-context instruction learning and collect a high-quality real dataset for comic book summarization,
@@ 3.1 Architecture
Stage 1 Stage 2 Stage 3 Stage 4 Sequence Length 32K 16K 128K 1M Batch Size 528 528 64 8 Training Iterations 1,000 5,000 1,000 500 Training Tokens 16B 40B 8B 4B Sequence Packing $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ Parallelism Tensor 8 8 8 8 Pipeline 1 1 1 1 Context 1 1 2 8 Data 8 8 4 1 Image Resolution $448+448\times\{1\times 2,2\times 1,\dots,3\times 4,4\times 3\}$ Max Number 128 64 512 4,096 Video Resolution $448$ FPS $1$ Max Frames 128 64 512 4,096 Learing Rate Vision $0.0$ $1.0\times 10^{-6}$ $1.0\times 10^{-6}$ $1.0\times 10^{-6}$ Projector $1.0\times 10^{-3}$ $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ LLM $0.0$ $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ Learing Rate Decay Vision $0.0$ $0.9$ Learing Rate scheduler Cosine Weight Decay $0.0$ Gradient Clip $1.0$ Rotary Base $1,000,000$ Adam Beta1 $0.9$ Adam Beta2 $0.999$
@@ 3.2 Data Construction
  • Interleaved Image-Text. To empower all-round multi-image capabilities, we employ the M4-Instruct [38]. To further enhance multi-image understanding with more than $10$ images, we collect the public comic book with the corresponding detailed synopsis from the web and build the Comic-9k datasets. Specifically, Comic-9k contains $200$K images, spanning $9$K comic books, along with a manual-labeled synopsis.
@@
Note that Comic-9k and MovieNet-Summary are created by this work and are made publicly available. Therefore, Long-VITA is only trained on open data, and we do not use data filtering methods.
@@ 3.3 Training Pipelines
Stage 2: General Knowledge Learning. After establishing the vision-language alignment in the embedding space, we dedicate most of our computational resources to vision-language general knowledge learning. This stage leverages all the image-text data for multiple tasks, including image captioning, common VQA, OCR, and multi-model conversations. In this stage, we also add text-only general instructions, math problems, and arithmetic calculations. For video understanding, we only add VideoGPT-plus [39] and ShareGemini-cap [40]. In both Stage 1 and Stage 2, we pack all training data to a fixed sequence length, which effectively trains samples with different lengths of sequences. Specifically, we random sample data items from the same source and concatenate them into one training sample with a token length of $32$K and $16$K for Stage 1 and Stage 2, respectively. We reset positional embeddings and attention masks for all packed samples so that each text-vision pair only attends to itself. This approach helps manage extensive datasets and ensure diverse data segments’ coverage.
Stage 3: Long-Sequence Fine-Tuning. In this stage, we extend the context length to $128$K. We reduce the sampling ratio of the data in Stage 2 to $0.1$ and incorporate additional long-context text instructions, comic book summaries, and video understanding datasets.
Stage 4: Long-Sequence Fine-Tuning. In this stage, we extend the context length to $1,024$K and add additional movie summary data. In both Stage 3 and Stage 4, we also pack all training data to a fixed sequence length without resetting positional embedding and attention mask. Therefore, we impose the model to capture the correlations between these two modalities in long-contextual information.
We do not use the interpolation technique during training and testing, therefore, the context window of Long-VITA can be extended further when equipped with YaRN [24], LongRoPE [25] and NTK-based interpolation. Note that we do not use any parameter-efficient methods such as LoRA [60] or approximate attention [26].
@@ 3.4 Hyper-parameters and Infrastructures
We initially implement training and inference with MindSpeed [78] and MindSpeed-LLM [79], which adapt Megatron-LM [80] to Ascend NPU. We also transfer the training and inference code to the GPU platform. As shown in Tab. 2, we list the detailed hyper-parameters in Long-VITA.
@@
  • Logits-Masked Language Modeling Head. We observe that the output logits from the language modeling head induce excessive memory footprints. For example, given $1$M tokens with $10^{5}$ vocabularies, the output logit matrix has a shape of $10^{6}\times 10^{5}$ and requires $400$ GB of memory for the float32 data type. To address the memory issue, we mask out all hidden features and only feed the one hidden feature that predicts the next tokens to the language modeling head. With the above design, the memory consumption of the output logit matrix is $0.0004$ GB with $10^{6}\times$ reduction. Note that this design can also apply to model training with long-context inputs and the language modeling head only needs to predict the short-context outputs to reduce memory consumption.
We test the maximal sequence length of a fixed number of devices before they raise the out-of-memory error. Tab. 3 summarizes the result. Note that activation checkpointing is disabled in LongVILA [11], while our model has much more number of parameters.
Tab. 4 further shows the effectiveness of logits-masked language modeling head (logits-masked LM head). All methods are implement with Flash Attention and context-parallelism distributed inference on GPU with $96$G memory and about $150$ TFLOPS for bfloat16. Compared to the original LM head, logits-masked LM head extends the max sequence length by $417\%$, and reduces time cost by $47.3\%$. We also implement a chunked language modeling head (chunked LM head), which process tokens with a chunk length of $32,768$. Compared to the chunked LM head, logits-masked LM head achieves $11.3$ and $11.1$ speedup under the $1$M and $1.6$M input lengths, respectively.
We employ Ring Attention [81] to distribute long sequences across multiple devices. We do not use parameter-efficient fine-tuning methods or quantization strategies for both training and inference. Additionally, the temperature is set to 00 to guarantee consistent performance evaluation.
@@ 4 Experiments
We evaluate Long-VITA ’s performance on image and video understanding with different sequence lengths, respectively. We perform a comprehensive evaluation on the OpenCompass benchmark, which covers visual question answering, multimodal conversation, knowledge and reasoning, OCR, and hallucination. OpenCompass is a comprehensive collection with $8$ popular multimodal benchmarks, including MMBench [20], MMStar [22], MMMU [82], MathVista [61], HallusionBench [62], AI2D [83], OCRBench [63], and MMVet [84]. We report the average scores for all collections. We also calculate the average scores of the objective benchmark only. We adopt Video-MME [29] as the video understanding evaluation indicator. Video-MME is an ideal benchmark for assessing LMMs’ ability to handle long videos in real-world scenarios, given its average video duration of 1017 seconds and the inclusion of short, medium, and long subsets. We use greedy search to ensure reproducible experimental results.
@@
  • Fully open models. Models are fully released with weights, training data, code, and evaluation, and thus can be fully reproduced.
@@ 4.2 Image Evaluation
Tab. 5 presents the main results on the OpenCompass leaderboard, which evaluates our model on various image benchmarks to investigate the image performance. As shown in Tab. 5, Long-VITA-16K demonstrates superior performance among open-source models. In most benchmarks, Long-VITA-16K surpasses Qwen2-VL-7B and InternVL2-8B. This highlights Long-VITA-16K’s impressive capabilities in handling multi-image tasks. Long-VITA-16K also achieves new state-of-the-art performance on MMMU and HallusionBench and outperforms strong open-source models by a notable margin. However, the results in Long-VITA-1M fall short of Long-VITA-16K and Long-VITA-128K, since we pack samples without isolating them via attention masks during 1M training and the potential confusion may arise from different sources of training data. In summary, Long-VITA achieves very competitive performance compared to other models with in-house data under the 20B parameters. This demonstrates that pure open-source data can also build robust models with strong performance.
@@ 4.3 Video Evaluation
Long-VITA models also show strong video understanding capabilities. Tab. 6 compares different models with various frame numbers on Video-MME. In particular, the effectiveness of Long-VITA is further underscored by its performance on Video-MME. Specifically, Long-VITA-128K with $256$ frames exceeds all other models under the 20B parameters. It shows exceptional results, especially in tasks involving medium- to long-length videos, outperforming other fully open video models such as LLaVA-Video-7B-Qwen [28] and LongVILA-7B [11]. Long-VITA-1M supports a maximum number of $4,096$ frames as input and achieves competitive performance. Note that Long-VITA is fully compatible with slow-fast [15] and progressive pooling [85] strategies, which are train-free video representations to extend the visual context window. Meanwhile, we do not adjust the scale factor of rotary position embedding during the pre-training and fine-tuning stages; thus employing existing interpretation methods [23, 24] can further achieve extrapolation during the inference phase.
To further demonstrate the exceptional long- and short-context understanding capability of our method, we conduct experiments on benchmarks on LongVideoBench [31] and MVBench [86] for long- and short-context video understanding tasks. The results are shown in Tab. 7, highlighting our model’s efficacy across varying temporal lengths. The Long-VITA-128K model is remarkably capable of understanding long-form video content. Specifically, our Long-VITA-128K model surpasses all existing $7$B to $20$B model series on LongVideoBench. Furthermore, the Long-VITA-1M model showcases strong performance across video understanding of $64$ to $4,096$ frames.
Remarkably, despite achieving these impressive results, Long-VITA is only learned from open-source data compared to other models. These results reflect a significant advancement in the efforts of the research community to close the performance gap with commercial models.
@@ 5 Conclusion
Contributions. This work introduces the Long-VITA series models as a primary exploration into powerful long-context vision-language models. Thanks to open-source data and training infrastructures, Long-VITA has excellent results compared to the cutting-edge models under $20$B parameters with in-house data and achieves new state-of-the-art performance for both image and video understanding on serval benchmarks.

### v3

@@ Abstract
[…]
@@ 1 Introduction
To further push the limits of open-source model capabilities, we extend the context length to $1$ million tokens and introduce Long-VITA, a strong open-source long-context visual language model. To this end, we employ a phased training approach that positions language as the pivot. Specifically, Long-VITA ’s ability to handle extended contexts is systematically augmented through a four-stage process. Beyond the conventional stages of vision-language alignment and supervised fine-tuning, our approach incorporates specialized stages for long-context supervised fine-tuning of 128128K and 11M. To achieve a good performance trade-off between long and short sequences, Long-VITA takes advantage of the existing abundance of open-source image-text and video-text data. We also introduce a multi-image summarization dataset, Comic-9K, comprising 99k comic books and the corresponding detailed synopsis. This dataset has a total of 200200k images with an average of $20$ high-resolution photos for each sample, and the synopsis is all manually written and collected from the web.
@@
  • We evaluate Long-VITA on a wide range of benchmarks. Although Long-VITA is trained only on open-source data, comprehensive evaluations reveal that Long-VITA has emerged as a strong vision-language model among previous models of similar scale, especially in evaluating hallucinations and video understanding.
@@ 2.2 Long-Context Multi-Modal Model
Many methods have also been proposed in large multi-modal models to handle long-context visual inputs. LongVILA [11] executes a continuation of pre-training on the LLM to enhance its context length to 256256K, followed by long video training. LongLLaVA [12] integrates a hybrid of Mamba and Transformer blocks for long-context multi-modal understanding. LongVU [13] proposes a spatio-temporal adaptive compression scheme to reduce long video tokens by leveraging cross-modal query and inter-frame similarities. LongVA [14] extends the language model on text data and then aligns the extended model with visual inputs, which perceives more than 200200K visual tokens. Kangaroo [27] develops a curriculum training strategy that progressively equips the LLM with the capacity to comprehend long videos.
@@ 2.3 Long-Context Visual Instruction Data
LLaVA-Video [28] synthesizes long video-language instruction data, covering various tasks such as captioning, open-ended, and multi-choice QA. LongVILA [11] constructs instruction-following datasets from long videos, which encompass summarization and other queries relevant to a comprehensive understanding of the content of long videos. Video-MME [29] incorporates a diverse range of video types, varying temporal durations, ranging from $11$ seconds to $1$ hour. LVBench [30] evaluates long-context LMMs that feature video-language interleaved inputs up to an hour long. LongVideoBench [31] introduces referring reasoning questions and presents significant challenges for both proprietary and open-source LMMs in their long-context multi-modal capabilities.
However, the above works only focus on long-context understanding in videos, which usually contain redundant visual frames and other modal information, such as subtitles and audio. In this paper, we explore the comic-based long-context instruction learning and collect a high-quality real dataset for comic book summarization.
@@ 3.1 Architecture
Stage 1 Stage 2 Stage 3 Stage 4 Sequence Length 32K 16K 128K 1M Batch Size 528 528 64 8 Training Iterations 1,000 5,000 1,000 500 Training Tokens 16B 40B 8B 4B Sequence Packing $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ Parallelism Tensor 8 8 8 8 Pipeline 1 1 1 1 Context 1 1 2 8 Data 8 8 4 1 Image Resolution $448+448\times\{1\times 2,2\times 1,\dots,3\times 4,4\times 3\}$ Max Number 128 64 512 4,096 Video Resolution $448$ FPS $1$ Max Frames 128 64 512 4,096 Learing Rate Vision $0.0$ $\times 10^{-6} 1.0×10$$\times 10^{-6}$ $1.0\times 10^{-6}$ Projector $\times 10^{-3} 1.0×10$$\times 10^{-5}$ $\times 10^{-5} 1.0×10$$\times 10^{-5}$ LLM $0.0$ $\times 10^{-5} 1.0×10$$\times 10^{-5}$ $1.0\times 10^{-5}$ Learing Rate Decay Vision $0.0$ $0.9$ Learing Rate scheduler Cosine Weight Decay $0.0$ Gradient Clip $1.0$ Rotary Base $1,000,000$ Adam Beta1 $0.9$ Adam Beta2 $0.999$
@@ 3.2 Data Construction
  • Interleaved Image-Text. To empower all-round multi-image capabilities, we employ the M4-Instruct [38]. To further enhance multi-image understanding with more than $10$ images, we collect the public comic book with the corresponding detailed synopsis from the web and build the Comic-9k datasets. Specifically, Comic-9k contains $200,000$ images, spanning $9,000$ comic books, along with a manually labeled synopsis.
@@
Note that Comic-9k and MovieNet-Summary were created by this work and are made publicly available. Therefore, Long-VITA is only trained on open data, and we do not use data filtering methods.
@@ 3.3 Training Pipelines
Stage 2: Learning of general knowledge. After establishing vision-language alignment in the embedding space, we dedicate most of our computational resources to general knowledge learning in the vision-language. This stage leverages all the image-text data for multiple tasks, including image captioning, common VQA, OCR, and multi-model conversations. In this stage, we also add text-only general instructions, math problems, and arithmetic calculations. For video understanding, we only add VideoGPT-plus [39] and ShareGemini-cap [40]. In both Stage 1 and Stage 2, we pack all training data to a fixed sequence length, which effectively trains samples with different lengths of sequences. Specifically, we randomly sample data items from the same source and concatenate them into one training sample with a token length of 3232K and 1616K for Stage 1 and Stage 2, respectively. We reset positional embeddings and attention masks for all packed samples so that each text-vision pair only attends to itself. This approach helps manage extensive datasets and ensure the coverage of diverse data segments.
Stage 3: Long-Sequence Fine-Tuning. In this stage, we extend the context length to 128128K. We reduce the sampling ratio of the data in Stage 2 to $0.1$ and incorporate additional long-context text instructions, comic book summaries, and video understanding datasets.
Stage 4: Long-Sequence Fine-Tuning. In this stage, we extend the context length to 1,0241,024K and add additional movie summary data. In both Stage 3 and Stage 4, we also pack all training data to a fixed sequence length without resetting positional embedding and attention mask. Therefore, we impose the model to capture the correlations between these two modalities in long-contextual information.
We do not use the interpolation technique during training and testing; therefore, the context window of Long-VITA can be extended further when equipped with YaRN [24], LongRoPE [25], and NTK-based interpolation. Note that we do not use any parameter-efficient methods such as LoRA [60] or approximate attention [26].
@@ 3.4 Hyper-parameters and Infrastructures
As shown in Tab. 2, we list the detailed hyperparameters in Long-VITA.
@@
  • Logits-Masked Language Modeling Head. We observe that the output logits from the language modeling head induce excessive memory footprints. For example, given 11M tokens with $10^{5}$ vocabularies, the output logit matrix has a shape of $10^{6}\times 10^{5}$ and requires $400$ GB of memory for the float32 data type. To address the memory issue, we mask out all hidden features and only feed the one hidden feature that predicts the next tokens to the language modeling head. With the above design, the memory consumption of the output logit matrix is $0.0004$ GB with $10^{6}\times$ reduction. Note that this design can also apply to model training with long-context inputs and the language modeling head only needs to predict the short-context outputs to reduce memory consumption.
We test the maximal sequence length of a fixed number of devices before they raise the out-of-memory error. Tab. 3 summarizes the result. Note that activation checkpointing is disabled in LongVILA [11], while our model has a much greater number of parameters.
Tab. 4 further shows the effectiveness of logits-masked language modeling head (logits-masked LM head). All methods are implemented with Flash Attention and context-parallelism distributed inference on GPU with 9696G memory and about $150$ TFLOPS for bfloat16. Compared to the original LM head, the logits-masked LM head extends the max sequence length by $417\%$, and reduces time cost by $47.3\%$. We also implement a chunked language modeling head (chunked LM head), which processes tokens with a chunk length of $32,768$. Compared to the chunked LM head, the logits-masked LM head achieves $11.3$ and $11.1$ speedup under the 11M and 1.61.6M input lengths, respectively.
We employ Ring Attention [78] to distribute long sequences across multiple devices. We do not use parameter-efficient fine-tuning methods or quantization strategies for both training and inference. Additionally, the temperature is set to 0 to guarantee consistent performance evaluation.
@@ 4 Experiments
We evaluate Long-VITA ’s performance on image and video understanding with different sequence lengths, respectively. We perform a comprehensive evaluation on the OpenCompass benchmark, which covers visual question answering, multimodal conversation, knowledge and reasoning, OCR, and hallucination. OpenCompass is a comprehensive collection with $8$ popular multimodal benchmarks, including MMBench [20], MMStar [22], MMMU [79], MathVista [61], HallusionBench [62], AI2D [80], OCRBench [63], and MMVet [81]. We report the average scores for all collections. We also calculate the average scores of the objective benchmark only. We adopt Video-MME [29] as the video understanding evaluation indicator. Video-MME is an ideal benchmark for evaluating LMM’s ability to handle long videos in real-world scenarios, given its average video duration of 1017 seconds and the inclusion of short, medium, and long subsets. We use greedy search to ensure reproducible experimental results.
@@
  • Fully open models. The models are fully released with weights, training data, code, and evaluation, and thus can be fully reproduced.
@@ 4.2 Image Evaluation
Tab. 5 presents the main results on the OpenCompass leaderboard, which evaluates our model on various image benchmarks to investigate image performance. As shown in Tab. 5, Long-VITA-16K demonstrates superior performance among open-source models. In most benchmarks, Long-VITA-16K exceeds Qwen2-VL-7B and InternVL2-8B. This highlights Long-VITA-16K’s impressive capabilities in handling multi-image tasks. Long-VITA-16K also achieves new state-of-the-art performance on MMMU and HallusionBench and outperforms strong open-source models by a notable margin. However, the results in Long-VITA-1M fall short of Long-VITA-16K and Long-VITA-128K, since we pack samples without isolating them via attention masks during 1M training, and the potential confusion may arise from different sources of training data. In summary, Long-VITA achieves very competitive performance compared to other models with in-house data under the 20B parameters. This demonstrates that pure open-source data can also build robust models with strong performance.
@@ 4.3 Video Evaluation
Long-VITA models also show strong video understanding capabilities. Tab. 6 compares different models with various frame numbers on Video-MME. In particular, the effectiveness of Long-VITA is further underscored by its performance on Video-MME. Specifically, Long-VITA-128K with $256$ frames exceeds all other models under the 20B parameters. It shows exceptional results, especially in tasks involving medium- to long-length videos, outperforming other fully open video models such as LLaVA-Video-7B-Qwen [28] and LongVILA-7B [11]. Long-VITA-1M supports a maximum number of $4,096$ frames as input and achieves competitive performance. Note that Long-VITA is fully compatible with slow-fast [15] and progressive pooling [82] strategies, which are train-free video representations to extend the visual context window. Meanwhile, we do not adjust the scale factor of rotary position embedding during the pre-training and fine-tuning stages; thus, employing existing interpretation methods [23, 24] can further achieve extrapolation during the inference phase.
To further demonstrate the exceptional long- and short-context understanding capability of our method, we perform benchmark experiments on LongVideoBench [31] and MVBench [83] for long- and short-context video understanding tasks. The results are shown in Tab. 7, highlighting our model’s efficacy across varying temporal lengths. The Long-VITA-128K model is remarkably capable of understanding long-form video content. Specifically, our Long-VITA-128K model surpasses all existing 77B to 2020B model series on LongVideoBench. Furthermore, the Long-VITA-1M model displays strong performance in video understanding of $64$ to $4,096$ frames.
Remarkably, despite achieving these impressive results, Long-VITA is only learned from open-source data compared to other models. These results reflect a significant advancement in the research community’s efforts to close the performance gap with commercial models.
@@ 5 Conclusion
Contributions. This work introduces the Long-VITA series models as a primary exploration into powerful long-context vision-language models. Thanks to open-source data and training infrastructures, Long-VITA has excellent results compared to the cutting-edge models under 2020B parameters with in-house data and achieves new state-of-the-art performance for both image and video understanding on several benchmarks.